Grade 11Mechanics

A cyclist completes half of a journey with speed V1 and the other with speed V2. Prove that average speed is 2V1V2/V1+V2

Vidura
9 Years agoGrade 11

2 Answers

Khimraj
ApprovedApproved Tutor Answer9 Years ago
Average speed is (total distance)/(time take). Let total distance is d then avg speed = d/(d/2V1) +(d/2V2). So avg speed is 2V1V2/V1+V2. Hope it clears. If u have doubts then please clearify. And if u like my answer then please approve answer.
manoj kumar reddy
8 Years ago
let us first consider that distance covered in first half be ‘x’ in T1 seconds, then the second half is also ‘x’ in T2 seconds.
since we know Average speed =(total distance)/(time take).
total distance=2x.............................................(1)
since TIME = (DISTANCE) x (SPEED)
then T1 = X / V1
and  T2 = X/V2
total time taken =T1+T2
= (X)/(V1) +(X)/(V2)
= X[1/V1+1/V2]
= X[(V1V2)(V1+V2)...................................(2)
AVEARGE SPEED =(TOTAL DISTANCE) / (TOTAL TIME TAKEN)
= (2X) / X(V1V2) / (V1+V2)
= ( 2) V1V2 / V1+ V2
I think this would help you